OpenStudy (anonymous):

Find the missing term of the arithmetic sequence 20, ___ , 32, ….

OpenStudy (anonymous):

32-20=12 12/2=6 Add 6 to the first number.

OpenStudy (whpalmer4):

or looking at it a slightly different way, if the common difference is \(d\), our sequence is \[20, 20+d, 32\]or\[20, 20+d, 20+d+d\]and \[20+d+d=32\] Easy enough to solve that for the value of \(d\), right?